OpenStudy (anonymous):

Well, I'm back with some more square roots. Simplify: 8√ x √5 x √2 x √20 The x's are multiplication signs.

OpenStudy (anonymous):

sqrt8 can be written as sqrt(4 x 2)

OpenStudy (anonymous):

\[\sqrt{8}=\sqrt{4*2}=2\sqrt{2}\]

OpenStudy (anonymous):

Oh, jeez.

OpenStudy (anonymous):

I'm gonna have to break them all down, huh?

OpenStudy (anonymous):

break down the sqrt 20 as well \[\sqrt{20}=\sqrt{4*5}=2\sqrt{5}\]

OpenStudy (anonymous):

no just the sqrt8 and sqrt20 the others can't be broken down any further

OpenStudy (anonymous):

P.S. you can do a √ by holding alt on your keyboard, then typing 251, then let go of Alt.

OpenStudy (anonymous):

(sqrt) is annoying. ._.

OpenStudy (anonymous):

Oh btw, type 251 on your numpad. not the ones above your keyboard. But anyways...

OpenStudy (anonymous):

I guess you can keep using sqrt. If you must. lol.

OpenStudy (anonymous):

dont have a numpad, its a small laptop...notebook thing

OpenStudy (anonymous):

Kay. Anyways, moving on then.

OpenStudy (anonymous):

anyway. can you figure it out from there ^^^

OpenStudy (anonymous):

Let me try real quick....

OpenStudy (anonymous):

Nope. I suck at simplifying square roots ;_;.

OpenStudy (anonymous):

oops!!! I misread the question

OpenStudy (anonymous):

just multiply everything together

OpenStudy (anonymous):

.-.

OpenStudy (anonymous):

then my original guess of √35 would be correct, yeah?

OpenStudy (anonymous):

that IS aa multiplication between each sqrt right?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

\[=\sqrt{8*2*5*20}\]\[\sqrt{8}*\sqrt{2}*\sqrt{5}*\sqrt{20}\]

OpenStudy (anonymous):

sorry, the 1st line was supposed to go underneath the 2nd one

OpenStudy (anonymous):

SO √35, right?

OpenStudy (anonymous):

That's the answer? It's one of my options. It's kinda an online quiz.

OpenStudy (anonymous):

That's what I'm doing.

OpenStudy (anonymous):

no, 8 x 2 x 5 x 20 = 1600 sqrt1600 = 40

OpenStudy (anonymous):

Wow. well then. √40 is also an option.

OpenStudy (anonymous):

So I'll go with that. Thanks....

OpenStudy (anonymous):

no, just 40

OpenStudy (anonymous):

Yeah. there's 40 and √40.

OpenStudy (anonymous):

just 40 :)

OpenStudy (anonymous):

Can't wait until square roots is over. I hate square roots now. Wasting hours of my time with this damned stuff.

OpenStudy (anonymous):

Very tricky and annoying.

OpenStudy (anonymous):

Well, I PERHAPS got the hang of square roots. I'm not even sure. Probably in another quiz I'll be thrown a problem that I get stuck on, and I'll probably have to come back to OpenStudy again to get help for it.... But anyways, thanks again. :/

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Square Roots seemed simple at first... I was wrong lol.

OpenStudy (anonymous):

well, ok bye :)